Operations Report – Feb 28th

Crew 222 Operations Report 28FEB2020

SOL: 12

Name of person filing report: Valentin PRUDHOMME

Non-nominal systems: Nothing to report

Notes on non-nominal systems: Nothing to report

Thursday generator:

Off at 5:30am

On at 9:00pm SOC was 83%

Friday generator:

Off at 5:00am

On at 9:00pm SOC was XX%

Diesel Reading – 70%

Station Propane Reading – 83%

Ethanol Free Gasoline – 7 gallons

Water (loft tank): 47 gallons

Water Meter: 149590.24 units

Water (static tank) – 176 gallons

Static to Loft Pump used – yes

Water in GreenHab – 0 gallons

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours: 146.7

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Spirit rover: no

Hours: 132.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: moved from campus to the shop to repair the brakes.

Hours: Nothing to report

Beginning charge: Nothing to report

Ending charge: Nothing to report

Currently charging: Nothing to report

Curiosity rover used: no

Hours: 138.8

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Notes on rovers: Nothing to report

ATV’s Used: yes (Honda, 1, 2 and 3)

Reason for use: Run during the engineer EVA for the battery and for the

EVA this morning.

Oil Added? no

ATV Fuel Used: 20% of the tank.

Hours the ATVs were Used today: 45 minutes

Notes on ATVs: Oil checked. The Honda and the ATV number 1’s oil light lit up. At the end of the EVA, the red one couldn’t change gear. I stopped and it was stuck in 5th gear. We had to push it to change gear. When we finally put it in neutral, we started it again and it worked well until the end (the gears were working fine).

HabCar used and why, where? The HabCar is now in the shop awaiting repairs.

CrewCar used and why, where? The CrewCar is in Grand Junction.

General notes and comments: Nothing to report

Summary of internet: All ok.

Summary of suits and radios: Suits 7, 8, 9, 10 were used during the EVA. The suit number 7 was working well at the beginning but stopped working in the middle of the EVA. We fixed this in the RAM during this afternoon and now it’s working well.

Summary of Hab operations: Nothing to report

Summary of GreenHab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Fixing of the suit number 7. It’s working well now. Finished the spacesuit experiment and all working well now.

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Nothing to report